OpenStudy (anonymous):

if acostheta-bsintheta=c then value of (asintheta+bcostheta) is a.(a^2+b^2-c^2)^1/2 b.(a^2-b^2+c^2)^1/2 c..(a^2+b^2+c^2)^1/2 d..(b^2+c^2-a^2)^1/2

OpenStudy (anonymous):

\[\begin{align*}a\sin\theta+b\cos\theta&=\sqrt{(a\sin\theta+b\cos\theta)^2}\\\\ &=\sqrt{a^2\sin^2\theta+2ab\sin\theta\cos\theta+b^2\cos^2\theta}\\\\ &=\sqrt{a^2\left(1-\cos^2\theta\right)+2ab\sin\theta\cos\theta+b^2\left(1-\sin^2\theta\right)}\\\\ &=\sqrt{a^2-a^2\cos^2\theta+2ab\sin\theta\cos\theta+b^2-b^2\sin^2\theta}\\\\ &=\sqrt{a^2+b^2-\left(a^2\cos^2\theta-2ab\sin\theta\cos\theta+b^2\sin^2\theta\right)}\\\\ &=\sqrt{a^2+b^2-\left(a\cos\theta-b\sin\theta\right)^2}\\\\ &=\sqrt{a^2+b^2-c^2}\\\\ \end{align*}\]

HanAkoSolo (jamierox4ev3r):

@Sandeepsra the porcess makes sense, I do not know haow to explain it any better!

HanAkoSolo (jamierox4ev3r):

oh and @SithsAndGiggles . nice job! couldn't have expalined it better myself, and nice usage of \(\LaTeX\) or the equation editor? idk but good job :)

OpenStudy (anonymous):

Thanks @Jamierox4ev3r !

OpenStudy (anonymous):

@Sandeepsra \[\begin{align*}a\sin\theta+b\cos\theta&=\sqrt{(a\sin\theta+b\cos\theta)^2}\\\\ \end{align*}\] This is like saying \(x=\sqrt{x^2}\). Make sense so far?

OpenStudy (anonymous):

In the next step, I expand the binomial: \[\begin{align*}(a\sin\theta+b\cos\theta)^2&=(a\sin\theta+b\cos\theta)(a\sin\theta+b\cos\theta)\\ &=a^2\sin^2\theta+2ab\sin\theta\cos\theta+b^2\cos^2\theta \end{align*}\] So our expression is the same as in the second line: \[\sqrt{a^2\sin^2\theta+2ab\sin\theta\cos\theta+b^2\cos^2\theta}\]

OpenStudy (anonymous):

Next, I use the Pythagorean identity, which says that \(\sin^2x=1-\cos^2x\) and \(\cos^2=1-\sin^2x\). Doing so gives me \[\sqrt{a^2\left(1-\cos^2\theta\right)+2ab\sin\theta\cos\theta+b^2\left(1-\sin^2\theta\right)}\]

OpenStudy (anonymous):

Then I distribute: \[a^2\left(1-\cos^2\theta\right)=a^2-a^2\cos^2\theta\] I do the same for \(b^2\left(1-\sin^2\theta\right)\). So I get \[\sqrt{a^2-a^2\cos^2\theta+2ab\sin\theta\cos\theta+b^2-b^2\sin^2\theta}\]

OpenStudy (anonymous):

The next step involves the commutative property of addition (\(a+b=b+a\)), where I simply move around some terms in the radicand: \[\sqrt{a^2-a^2\cos^2\theta+2ab\sin\theta\cos\theta+b^2-b^2\sin^2\theta}\\ ~~~~~~=\sqrt{a^2+b^2-a^2\cos^2\theta+2ab\sin\theta\cos\theta-b^2\sin^2\theta}\] Next, I factor out a -1 from the last three terms. \[\sqrt{a^2+b^2-\left(a^2\cos^2\theta-2ab\sin\theta\cos\theta+b^2\sin^2\theta\right) }\]

OpenStudy (anonymous):

This next part is a lot like the first. The quantity within parentheses can be factored as a squared binomial: \[\begin{align*}a^2\cos^2\theta-2ab\sin\theta\cos\theta+b^2\sin^2\theta&=(a\cos\theta-b\sin\theta)(a\cos\theta-b\sin\theta)\\ &=(a\cos\theta-b\sin\theta)^2\end{align*}\] So I get \[\sqrt{a^2+b^2-\left(a\cos\theta-b\sin\theta\right)^2}\]

OpenStudy (anonymous):

Finally, I use the given expression for \(c\). Since \(c=a\cos\theta-b\sin\theta\), you get \(\sqrt{a^2+b^2-c^2}\).

HanAkoSolo (jamierox4ev3r):

man if i had another medal i would give it to you, I am not even kidding @SithsAndGiggles , that was extremely through :P

OpenStudy (anonymous):

Thanks again. Asker was a bit confused about the steps

HanAkoSolo (jamierox4ev3r):

haha apparently, although i don't see how you could know, he is not saying anything at all!

OpenStudy (anonymous):

Sent some PMs